OpenStudy (anonymous):

What is the missing number in this list? 17, 22, 27, _?_, 37

OpenStudy (anonymous):

32

OpenStudy (anonymous):

aww thank you !!

OpenStudy (whpalmer4):

22-17 = 5 27-22 = 5 37-5 = x x - 5 = 27 x = 32